# Advancing Prompt Learning through an External Layer

### v1

@@ Abstract.
Prompt learning represents a promising method for adapting pre-trained vision-language models (VLMs) to various downstream tasks by learning a set of text embeddings. One challenge inherent to these methods is the poor generalization performance due to the invalidity of the learned text embeddings for unseen tasks. A straightforward approach to bridge this gap is to freeze the text embeddings in prompts, which results in a lack of capacity to adapt VLMs for downstream tasks. To address this dilemma, we propose a paradigm called EnPrompt with a novel External Layer (EnLa). Specifically, we propose textual external layer and learnable visual embeddings for adapting VLMs to downstream tasks. The learnable external layer is built upon valid embeddings of pre-trained CLIP. This design considers the balance of learning capabilities between the two branches. To align the textual and visual features, we propose a novel two-pronged approach: i) we introduce the optimal transport as the discrepancy metric to align the vision and text modalities, and ii) we introduce a novel strengthening feature to enhance the interaction between these two modalities. Four representative experiments (i.e., base-to-novel generalization, few-shot learning, cross-dataset generalization, domain shifts generalization) across 15 datasets demonstrate that our method outperforms existing prompt learning methods in a wide range of datasets.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.isbn: 979-8-4007-0686-8/24/10doi: 10.1145/3664647.3680953ccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Seminal explorations involve hand-crafted text prompts, e.g., ‘a photo of a [class]’ as the prompt for text encoder. Advanced works propose to introduce a set of learnable parameters to adapt VLMs to downstream tasks. For instance, CoOp (Zhou et al., 2022c) keeps the weights of CLIP frozen while learning the text embeddings of prompts for efficient task-specific adaptation. These prompt-learning approaches achieve significant performance improvements over manually tuned prompts on various typical scenarios (Chen et al., 2023b; Cao et al., 2024c; Etchegaray et al., 2024). However, the text embeddings learned typically produce worse performance compared with hand-crafted prompts (Radford et al., 2021) for unseen tasks. Specifically, applying these learned text prompts to unseen classes leads to degenerated model generalization performance. Recent works understand and mitigate this problem through the overfitting view. To overcome the challenge, an image-conditional prompt (CoCoOp) learning approach (Zhou et al., 2022b) is proposed to promote the generalization performance in unseen classes, which is higher than CoOp. Unfortunately, CoCoOp exhibits lower generalization performance compared to hand-crafted zero-shot CLIP when applied to novel classes, as depicted in Table 4. This could be attributed to the learned embeddings are usually invalid in novel classes.
To address this dilemma, we propose a paradigm called EnPrompt with a novel External Layer (EnLa). Specifically, we propose textual external layer and learnable visual embeddings for adapting VLMs to downstream tasks through freezing the text embeddings. To align the textual and visual features, we introduce the optimal transport as a discrepancy metric that measures the difference between the visual and textual features, where the optimal transport can calculate the distance between two distributions under the form of multiple sampling. Further, we connect the text and image encoders through a strengthening feature instead of ordinary coupling for the dimension transformation of learnable hidden vectors. As shown in Figure 1, EnPrompt outperforms the existing state-of-the-art method in base-to-novel generalization task. Our main contributions can be summarized as follows:
  • We address the performance-degeneration issue through introducing a novel External Layer (EnLa) and freezing the text embeddings.
@@ 2.1. Preliminaries
We provide a brief introduction to vision-language pre-training, with a specific focus on CLIP, which is a zero-shot learning approach without fine-tuning. We build our approach on the basis of CLIP. CLIP has a text encoder $\mathcal{F}_{t}(\cdot)$ and an image encoder $\mathcal{F}_{v}(\cdot)$, which separately map a textual input (a.k.a. prompt) $\mathbf{p}$ and a visual input, i.e., an image, $\boldsymbol{x}$ into a shared feature space through many transformer blocks. The outputs of two encoders are denoted as ${\boldsymbol{t}}=\mathcal{F}_{t}(\mathbf{p})$ and ${\boldsymbol{v}}=\mathcal{F}_{v}(\mathbf{x})$. The image encoder aims to transform the input images into feature embeddings, while the text encoder generates representations for word embedding sequences.
Figure 2. EnPrompt (Ours) aims to connect the image and text encoders by a novel designs, which is introducing the strengthening feature of the EnLa to fuse the learnable visual embeddings of the first layer of image encoder, while frozen text embeddings. Both CoOp and MaPLe employ designs to learn text embeddings. In our method, the fusion embeddings are further introduced in transformer of the image encoder for deep prompting.
@@
where $y\in\mathcal{Y}$ is the label of $\mathbf{x}\in\mathcal{X}$, $\mathcal{P}=\left\{\mathbf{p}_{i}\right\}_{i=1}^{C}$ denotes the set of $C$ pre-defined prompts, $sim(\cdot,\cdot)$ stands for cosine similarity between two vectors, and $\tau>0$ represents a temperature parameter. Here, the classifier consists of $C$ textual features derived from pre-defined prompts $\mathcal{P}=\left\{\mathbf{p}_{i}\right\}_{i=1}^{C}$, where the prompt $\mathbf{p}_{i}$ for the $i$-th class may have the form of ‘a photo of a dog’.
Advanced works take a further step to investigate the possibility of aligning images and prompts. The insight of these works is that the prompts tuning for given images could be superior to hand-crafted prompts. Specifically, the class name is retained as prior knowledge to ensure the learned prompts can form a classifier, while the word (a.k.a. context) embeddings of prompts are modeled as learnable parameters. Here, the learnable words in the above prompt are typically initialized using ‘a photo of a’. The embeddings optimization approach can be formalized as follows,
@@
where ⓒ is a concatenating operation. Here, some learned text embeddings methods follow CoOp (Zhou et al., 2022c) to set the embeddings $\mathbf{w}$ shared across classes. In the inference phase, the prompts with the learned embeddings can produce textual features for classification.
@@ 2.2. EnLa with Frozen Text Embeddings
EnLa Design Analysis. We compare the performance of a single layer with a two-layer, referring to the bottleneck structure (Linear-ReLU-Linear), on 11 datesets with 16 shots for base-to-novel tasks. HM refers to the harmonic mean. The results in Table 1 indicate that the single 512x512 structure (row-5) provides better performance. It achieves the highest HM of $80.64\%$, which surpasses all reduction factors of the hidden layers that have 32x, 16x, 8x, and 4x. Thus, we used a single layer of EnLa in all of our experiments.
Table 1. Ablation study on the EnLa designs.
Method Reduction factor Base Acc Novel Acc. HM
@@ 2.3. Alignment with Optimal Transport
Given two sets that contain $N$ and $M$ points respectively, and the discrete distributions can be formulated as follows:
@@
(7) $\quad\mathbf{T}\mathbf{1}_{M}=\boldsymbol{\theta},\ \mathbf{T}^{\top}\mathbf{1}_{N}=\boldsymbol{\beta},$
where <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes the Frobenius dot-product and $\mathbf{1}_{N}$is the $N$ dimensional vector of ones. $\mathbf{C}\in\mathbb{R}_{>0}^{N\times M}$ is the cost matrix of the transport, and $C_{nm}$ denotes the transport cost between points $\boldsymbol{e}_{n}$ and $\boldsymbol{l}_{m}$, such as the cosine distance $C_{nm}=1-\operatorname{cosine}\left(\boldsymbol{e}_{n},\boldsymbol{l}_{m}\right).\mathbf{T}\in\mathbb{R}_{>0}^{N\times M}$ denotes the transport plan to be learned. OT distance is then minimized over all the joint probabilities of $N\times M$ space with two marginal constraints. As computing the above OT distance has the cubic time complexity, we apply the Sinkhorn distance  (Cuturi, 2013) that regularizes with an entropic constraint:
@@
(9) with $\mathbf{T1}_{M}=\boldsymbol{\theta},\ \mathbf{T}^{\top}\mathbf{1}_{N}=\boldsymbol{\beta}$
@@
(10) $d_{\mathrm{OT}}(k)=d_{\mathrm{OT}}\left(\boldsymbol{\theta},\boldsymbol{\beta}\mid\mathbf{1}-\boldsymbol{V}^{\top}\boldsymbol{t}_{k}\right),$
where $\boldsymbol{C}=\mathbf{1}-\boldsymbol{V}^{\top}\boldsymbol{t}_{k}$ denotes that we use the cosine distance between $\boldsymbol{V}$ and $\boldsymbol{t}_{k}$ as the cost matrix. Then, we can obtain the solution of transport plan $\mathbf{T}^{*}$ and the final OT distance $d_{\mathrm{OT}}(k)$. Given the OT distance between $\boldsymbol{t}_{k}$ and $\boldsymbol{V}$, and image $\boldsymbol{x}$, we reformulate the final prediction probability of V-L alignment as follows:
@@ 2.4. Alignment with Strengthening Feature
We introduce learnable visual embeddings for enhance learnality of VLMs for downstream tasks. In contrast to VPT (Jia et al., 2022), our method is being explored in the multi-modal prompt designs. Specifically, let $\mathcal{T}_{\epsilon}(\cdot)$ denote the V-L function parameterized by $\epsilon$ to transfer features from text branch to visual branch. The $\mathcal{T}_{\epsilon}(\cdot)$ is a function net with dimensionality of [512, 768] for dimension alignment of the visual branch and text branch. The V-L function will modify visual embeddings using the features generated by the EnLa, the input feature of the V-L function is the output produced by the EnLa. Let $\boldsymbol{e}$ denote the output of EnLa, Further, we append learnable $L$ visual input embeddings $\mathcal{P}_{v}$, which is the vector with [$L$, 768], given as follows,
@@
Table 2. Different positions of strengthening feature in base-to-novel generalization. Text embeddings frozen based on EnLa is the default setting. Results are averaged over 11 datasets. HM refers to the harmon.
@@
The hidden layer of the image encoder (row-1) combines the benefits of prompting in both branches by enforcing the image encoder representation ability through text knowledge transfer to the deep layer of the image encoder without visual embeddings initialization layer, and the input layer of the image encoder (row-2) is based on the connection of the visual embeddings initialization. The position (row-2) is more effective than position (row-1) in unseen tasks. It can be attributed to that position (row-1) lacks visual embedding initialization, leading to the acquisition of a limited ability to adapt the image encoder with the unbalance of V-L alignment. As a consequence, in our subsequent comprehensive experiments, we adopted the connection position with the input layer of the image encoder (row-2) in our approach. To this end, let $\boldsymbol{p}_{\boldsymbol{v}}(\boldsymbol{e})$ denote the fusion embeddings, which can be formalized as follows,
@@
In contrast to EnPrompt (Ours), CoCoOp transfers the output features of the image encoder to the text embeddings, which is more inference time consumption. MaPLe initializes the multi-learnable hidden vectors for V-L alignment and couple the text embeddings to the visual encoder. The coupling function is mapping these initialized hidden vectors of deep layer and text embeddings with the image encoder and text encoder for more parameters. However, our approach introduces the learnable visual initialization embeddings of the image encoder with the strengthening feature fusion. However, MaPLe learns to visual branch with the coupled visual embeddings without the visual embeddings initialization step, which limits the ability of visual branch to adapt downstream data distributions.
@@ 3.1. Benchmark Setting
We compare the performance with CLIP (Radford et al., 2021), CoOp (Zhou et al., 2022c), Clip-Adapter (Gao et al., 2023), CoCoOp (Zhou et al., 2022b), PLOT (Chen et al., 2022a), and MaPLe (Khattak et al., 2023). Please note that CLIP was originally a zero-shot method. However, in this context, we linearly process it using a few-shot method. The Clip-Adapter (Gao et al., 2023) is a frozen prompt method with the adapter component behind the encoder, which is different from EnPrompt. We compared it with PLOT (Chen et al., 2022a), which is an optimal transport method. PLOT optimizes four sentences simultaneously (”a photo of a dog”, ”a picture of a dog”, ”a drawing of a dog”, ”a good drawing of a dog”). The output text feature of EnPrompt is global feature, the input sentence is one text prompt such as ”a photo of a”, and the text prompt is non-learnable.
@@ 3.2. Effect of Components
Table 3. Effect of our components in base-to-novel generalization.
@@
5: + Strengthening feature with connecting 84.71 76.90 80.64
@@ 4.2. Training Strategy
Table 8. Comparison of EnLa train strategies. Results are averaged over 11 datasets. HM refers to harmonic mean
@@ 4.4. Visual Learning Depth
4.5. Analyzing of Multi-modal Designs
To validate the effect of EnPrompt (Ours), we conducted the HM task and few-shot learning task using two-modal designs and the method completely opposite to EnPrompt in Table 9. We observed that the method (row-3) has a lower performance in novel classes than the method (row-4), which is due to the learnable text embeddings of method (row-3) being weak in generalization compared to the frozen text prompt. Finally, we observed that the method (row-4) and method (row-3) performed better in the novel classes than method (row-2) while still maintaining a few-shot learning task. This is attributed to the fact that method (row-2) is to have neither learnable visual embeddings nor learnable text embeddings. It indicates that prompt learning is crucial for adapting pre-trained VLMs for generalization.

### v3

@@ Abstract.
Prompt learning represents a promising method for adapting pre-trained vision-language models (VLMs) to various downstream tasks by learning a set of text embeddings. One challenge inherent to these methods is the poor generalization performance due to the invalidity of the learned text embeddings for unseen tasks. A straightforward approach to bridge this gap is to freeze the text embeddings in prompts, which results in a lack of capacity to adapt VLMs for downstream tasks. To address this dilemma, we propose a paradigm called EnPrompt with a novel External Layer (EnLa). Specifically, we propose a textual external layer and learnable visual embeddings for adapting VLMs to downstream tasks. The learnable external layer is built upon valid embeddings of pre-trained CLIP. This design considers the balance of learning capabilities between the two branches. To align the textual and visual features, we propose a novel two-pronged approach: i) we introduce the optimal transport as the discrepancy metric to align the vision and text modalities, and ii) we introduce a novel strengthening feature to enhance the interaction between these two modalities. Four representative experiments (i.e., base-to-novel generalization, few-shot learning, cross-dataset generalization, domain shifts generalization) across 15 datasets demonstrate that our method outperforms the existing prompt learning method.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.isbn: 979-8-4007-0686-8/24/10doi: 10.1145/3664647.3680953ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
@@ 1. Introduction
Seminal explorations involve hand-crafted text prompts, e.g., “a photo of a [class]” as the prompt for text encoder. Advanced works introducing a set of learnable parameters to adapt VLMs to downstream tasks. For instance, CoOp (Zhou et al., 2022c) keeps the weights of CLIP frozen while learning the text embeddings of prompts for efficient task-specific adaptation. These prompt-learning approaches achieve significant performance improvements over manually tuned prompts on various typical scenarios (Chen et al., 2023b; Cao et al., 2024c; Etchegaray et al., 2024). However, the learned text embeddings typically produce worse performance compared with hand-crafted prompts (Radford et al., 2021) for unseen tasks. Specifically, applying these learned text prompts to unseen classes leads to degenerated model generalization performance. Recent works mitigate this problem through the overfitting view. To overcome the challenge, an image-conditional prompt (CoCoOp) learning approach (Zhou et al., 2022b) is proposed to improve the generalization performance on unseen classes. Unfortunately, CoCoOp exhibits lower generalization performance than hand-crafted zero-shot CLIP when applied to novel classes, as depicted in Table 4. This could be attributed to the learned embeddings are usually invalid on novel classes.
To address this dilemma, we propose a paradigm called EnPrompt with a novel External Layer (EnLa). Specifically, we propose adding a textual external layer on top of frozen text embeddings and using learnable visual embeddings for adapting VLMs to downstream tasks. To align the textual and visual features, we introduce the optimal transport (Villani et al., 2009) as a discrepancy metric that measures the difference between the visual and textual features, where the optimal transport can calculate the distance between two distributions under the form of multiple sampling. Further, we connect the text and image encoders through a strengthening feature instead of ordinary coupling for the dimension transformation of learnable hidden vectors. As shown in Figure 1, EnPrompt outperforms the existing state-of-the-art method on the base-to-novel generalization task. Our main contributions can be summarized as follows:
  • We address the performance-degeneration issue of VLMs on unseen tasks by introducing a novel External Layer (EnLa) and freezing the text embeddings.
@@ 2.1. Preliminaries
We provide a brief introduction to vision-language pre-training, with a specific focus on CLIP, which is a zero-shot learning approach without fine-tuning. We build our approach based on CLIP. CLIP has a text encoder $\mathcal{F}_{t}(\cdot)$ and an image encoder $\mathcal{F}_{v}(\cdot)$, which separately map a textual input (a.k.a. prompt) $\mathbf{p}$ and a visual input, i.e., an image, $\boldsymbol{x}$ into a shared feature space through many transformer blocks. The outputs of two encoders are denoted as ${\boldsymbol{t}}=\mathcal{F}_{t}(\mathbf{p})$ and ${\boldsymbol{v}}=\mathcal{F}_{v}(\mathbf{x})$. The image encoder aims to transform the input images into feature embeddings, while the text encoder generates representations for word embedding sequences.
Figure 2. Comparison of our method with CoOp and MaPLe. Our method freezes the textual embeddings and learn an External Layer (EnLa) on top of it. Optimal transport is introduced to align and visual and textual modalities apart from the conventional cross entropy (CE) loss. The textual and visual encoders are connected through a strengthening feature. This feature and the learnable visual prompt are fused and utilized for deep prompting of the visual encoder.
@@
where $y\in\mathcal{Y}$ is the label of $\mathbf{x}\in\mathcal{X}$, $\mathcal{P}=\left\{\mathbf{p}_{i}\right\}_{i=1}^{C}$ denotes the set of $C$ pre-defined prompts, $sim(\cdot,\cdot)$ stands for cosine similarity between two vectors, and $\tau>0$ represents a temperature parameter. Here, the classifier consists of $C$ textual features derived from pre-defined prompts $\mathcal{P}=\left\{\mathbf{p}_{i}\right\}_{i=1}^{C}$, where the prompt $\mathbf{p}_{i}$ for the $i$-th class may have the form of “a photo of a [class]”.
Advanced works take a further step to investigate the possibility of aligning images and prompts. The insight of these works is that the prompts tuning for given images could be superior to hand-crafted prompts. Specifically, the class name is retained as prior knowledge to ensure the learned prompts can form a classifier, while the word (a.k.a. context) embeddings of prompts are modeled as learnable parameters. Here, the learnable words in the above prompt are typically initialized using “a photo of a []”. The embeddings optimization approach can be formalized as follows,
@@
where ⓒ is the concatenating operation. Follow CoOp (Zhou et al., 2022c), the embeddings $\mathbf{w}$ can be shared across classes. In the inference phase, the prompts with the learned embeddings can produce textual features for classification.
@@ 2.2. EnLa with Frozen Text Embeddings
EnLa Design Analysis. We compare the performance of a single layer with a two-layer, referring to the bottleneck structure (Linear-ReLU-Linear), on 11 datasets with 16 shots for base-to-novel tasks. HM refers to the harmonic mean (Xian et al., 2017). The results in Table 1 indicate that the single 512x512 structure (row-5) provides better performance. It achieves the highest HM of $80.64\%$, which surpasses all reduction factors of the hidden layers that have 32x, 16x, 8x, and 4x. Thus, we used a single layer of EnLa in all of our experiments.
Table 1. Ablation study on the different designs of EnLa. Results are averaged over 11 datasets.
Layer Design Reduction factor Base Acc Novel Acc. HM
@@ 2.3. Alignment with Optimal Transport
Given two sets that contain $N$ and $M$ points respectively, the discrete distributions can be formulated as follows:
@@
[…]
where <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes the Frobenius dot-product and $\mathbf{1}_{N}$ is the $N$ dimensional vector of ones. $\mathbf{C}\in\mathbb{R}_{>0}^{N\times M}$ is the cost matrix of the transport, and $C_{nm}$ denotes the transport cost between points $\boldsymbol{e}_{n}$ and $\boldsymbol{l}_{m}$, such as the cosine distance $C_{nm}=1-\operatorname{cos}\left(\boldsymbol{e}_{n},\boldsymbol{l}_{m}\right).\mathbf{T}\in\mathbb{R}_{>0}^{N\times M}$ denotes the transport plan to be learned. OT distance is then minimized over all the joint probabilities of $N\times M$ space with two marginal constraints. As computing the above OT distance has the cubic time complexity, we apply the Sinkhorn distance  (Cuturi, 2013) that regularizes with an entropic constraint:
@@
[…]
@@
(10) $d_{\mathrm{OT}}(k)=d_{\mathrm{OT}}\left(\boldsymbol{\theta},\boldsymbol{\beta}\mid\mathbf{1}-\boldsymbol{V}^{\intercal}\boldsymbol{t}_{k}\right),$
where $\boldsymbol{C}=\mathbf{1}-\boldsymbol{V}^{\intercal}\boldsymbol{t}_{k}$ denotes that we use the cosine distance between $\boldsymbol{V}$ and $\boldsymbol{t}_{k}$ as the cost matrix. Then, we can obtain the solution of transport plan $\mathbf{T}^{*}$ and the final OT distance $d_{\mathrm{OT}}(k)$. Given the OT distance between $\boldsymbol{t}_{k}$ and $\boldsymbol{V}$, and image $\boldsymbol{x}$, we reformulate the final prediction probability of V-L alignment as follows:
@@ 2.4. Alignment with Strengthening Feature
We introduce learnable visual embeddings for enhancing learnability of VLMs for downstream tasks. In contrast to VPT (Jia et al., 2022), our method employ a multi-modal prompt design. Specifically, let $\mathcal{T}_{\epsilon}(\cdot)$ denote the V-L function parameterized by $\epsilon$ to transfer features from textual branch to visual branch. The $\mathcal{T}_{\epsilon}(\cdot)$ is a function net with dimensionality of [512, 768] for dimension alignment of the visual branch and text branch. The input feature of the V-L function is the output produced by the EnLa. Further, we append $L$ learnable visual input embeddings $\mathcal{P}_{v}$, which is a vector with [$L$, 768], given as follows,
@@
Table 2. Comparisons of different positions of strengthening feature in base-to-novel generalization. Results are averaged over 11 datasets.
@@
The hidden layer of the image encoder (row-1) combines the benefits of prompting in both branches by enforcing the image encoder representation ability through text knowledge transfer to the deep layer of the image encoder without visual embeddings initialization layer, and the input layer of the image encoder (row-2) is based on the connection of the visual embeddings initialization. The position (row-2) is more effective than the position (row-1) in unseen tasks. It can be attributed to that position (row-1) lacks visual embedding initialization, leading to a limited ability to adapt the image encoder with the unbalance of V-L alignment. As a consequence, in our subsequent comprehensive experiments, we adopted the connection position with the input layer of the image encoder (row-2) in our approach. Let $\boldsymbol{e}$ denote the output of EnLa, and let $\boldsymbol{p}_{\boldsymbol{v}}(\boldsymbol{e})$ denote the fusion embeddings, it can be expressed as:
@@
In contrast to EnPrompt (Ours), CoCoOp transfers the output features of the image encoder to the text embeddings, which is more inference time consumption. MaPLe initializes the multi-learnable hidden vectors for V-L alignment and couples the text embeddings to the visual encoder. The coupling function maps these initialized hidden vectors of deep layer and text embeddings with the image encoder and text encoder for more parameters. However, our approach introduces the learnable visual initialization embeddings of the image encoder with the strengthening feature fusion. However, MaPLe learns to visual branch with the coupled visual embeddings without the visual embeddings initialization step, which limits the ability of the visual branch to adapt downstream data distributions.
@@ 3.1. Benchmark Setting
We compare the performance with CLIP (Radford et al., 2021), CoOp (Zhou et al., 2022c), Clip-Adapter (Gao et al., 2023), CoCoOp (Zhou et al., 2022b), PLOT (Chen et al., 2022a), and MaPLe (Khattak et al., 2023). Please note that CLIP was originally a zero-shot method. However, in this context, we linearly process it using a few-shot method. The Clip-Adapter (Gao et al., 2023) is a frozen prompt method with the adapter component behind the encoder, which is different from EnPrompt. We compared it with PLOT (Chen et al., 2022a), which is an optimal transport method. PLOT optimizes four sentences simultaneously (“a photo of a dog”, “a picture of a dog”, “a drawing of a dog”, “a good drawing of a dog”). The output text feature of EnPrompt is global feature, the input sentence is one text prompt such as “a photo of a”, and the text prompt is non-learnable.
@@ 3.2. Effect of Components
Table 3. Effect of our components in base-to-novel generalization. Results are averaged over 11 datasets.
@@
5: + Strengthening feature 84.71 76.90 80.64
@@ 4.2. Training Strategy
Table 8. Comparison of EnLa train strategies. Results are averaged over 11 datasets.
@@ 4.4. Visual Learning Depth
4.5. Analyzing of Multi-modal Methods
To validate the effect of EnPrompt (Ours), we conducted the HM task and few-shot learning task using two-modal designs and the method completely opposite to EnPrompt in Table 9. We observed that the method (row-2) has a lower performance in novel classes than the method (row-3), which is due to the learnable text embeddings of method (row-2) being weak in generalization compared to the frozen text prompt. Finally, we observed that the method (row-3) and method (row-2) performed better in the novel classes than method (row-1) while still maintaining a non-generalization few-shot learning task. This is attributed to the fact that method (row-1) is to have neither learnable visual embeddings nor learnable text embeddings. It indicates that prompt learning is crucial for adapting pre-trained VLMs for generalization.